OpenStudy (anonymous):

Find the area of the shaded portion in the equilateral triangle with sides 6 assuming the central point of each arc is its corresponding vertex.

OpenStudy (anonymous):

do you have a photo

OpenStudy (anonymous):

Angle of arc = 60 deg Area of shaded region = area of triangle - 3(area of arc) r = 3 b = 6 h^2 + 3^2 = 6^2 h^2 = 36 - 9 h^2 = 27 h = 3√3 Area of triangle: A = (1/2)(6)(3√3) = 3(3√3) = 9√3 Area of arc: r = 3 angle = 60 deg Area of arc = (1/6)(area of circle) Area of circle: A = pi (3^2) = 9pi Area of arc = (1/6)(9pi) = 3pi/2 There are 3 such arcs: Area of arcs = 3(3pi/2) = 9pi/2 Area of shaded region: 9√3 - 9pi/2 = 1.45

OpenStudy (anonymous):

thank you so much!

OpenStudy (anonymous):

u r welcome :)